If 60% of the voters in a large district prefer candidate A, What is probability that in a sample of 12 voters exactly 7 will prefer A.

can you explain this question please.?


1
Expert's answer
2021-08-23T16:14:01-0400

Let "X=" the number of voters in a sample of 12 voters who will prefer A: "X\sim Bin(n, p)."

Given "n=12, p=0.6, q=1-p=0.4."


"P(X=7)=\dbinom{12}{7}(0.6)^7(0.4)^{12-7}"

"=792(0.6)^7(0.4)^{5}=0.227030335488\approx0.22703"